# A Game-Theoretic Unital Classification Theorem for $C^{*}$-Algebras

### v1

@@ 2.5. Borel category of $L$-structures
The map $\pi$ is precisely the quotient map under the equivalence $(A,B,f)\sim(A,B,g)\iff\widehat{f}=\widehat{g}$, and the above lemma verifies that the quotient is itself a standard Borel space, and $\pi$ is a Borel map. In the language of Borel equivalence relations, this means that $\sim$ is smooth (see for example [Kechris_2024, Definition 3.11]). We have successfully isolated a standard Borel space – namely $\pi(X_{1})$ – that is in natural bijection with the family of morphisms in $\mathcal{C}_{T}$, and will serve as the space of morphisms in the Borel category $\mathbf{X}^{T}$. Thus, we write $X^{T}_{\operatorname{Mor}}:=\pi(X_{1})$. We define $s,t\colon X^{T}_{\operatorname{Mor}}\to X^{T}_{\operatorname{Ob}}$ be the first and second projections respectively, and let $i\colon X^{T}_{\operatorname{Ob}}\to X^{T}_{\operatorname{Mor}}$ be the map sending $A$ to $\pi(A,A,i_{A})$, where $i_{A}(a,b)=d_{A}(a,b)$, i.e., is the identity map on $A$. We also define $\circ\colon X^{T}_{\operatorname{Mor}}\times_{t,X^{T}_{\operatorname{Ob}},s}X^{T}_{\operatorname{Mor}}\to X^{T}_{\operatorname{Mor}}$ to come from composition of homomorphisms in $\mathcal{C}_{T}$.
@@
The Borel category $\operatorname{\bf{Mod}}_{\omega}(T)$ is a Borel model of $\mathcal{C}_{T}$.
@@
Let $\widehat{\,\cdot\,}\colon\operatorname{Mod}_{\omega}(T)\to\mathcal{C}_{T}$ take any $A\in X^{T}_{\operatorname{Ob}}$ to $\widehat{A}$ and any $(A,B,F)\in X^{T}_{\operatorname{Mor}}(A,B)$ to the unique L{L}-homomorphism $\widehat{f}\colon\widehat{A}\to\widehat{B}$ that induces $F$. We have constructed $s,t,i$ and $\circ$ precisely such that this is a functor. Moreover, for fixed $A,B\in X^{T}_{\operatorname{Ob}}$, $\widehat{\,\cdot\,}$ is bijective from $X^{T}_{\operatorname{Mor}}(A,B)$ to the set of $L$-homomorphisms by construction (see Lemma 2.37). In other words, $\widehat{\,\cdot\,}$ is fully faithful.
@@ 3. Application to $C^{*}$-algebras
The desire to classify objects is ubiquitous across mathematical disciplines; here we discuss one particular classification program for norm-closed subalgebras of bounded operators of Hilbert spaces, i.e., $C^{*}$-algebras. See [CGSTW-new-classificationI-2023, § 1.1, § 1.2] for a survey and brief history of this topic. We assume familiarity with $C^{*}$-algebras, whose model theory is laid out in [farah2014model, § 2.3.1], see also [farah2021MTofC*]. We start by presenting the unital classification theorem along with the classifying invariant, in order to set out a first-order language suitable to understanding the invariant.
@@ 3.1. Borel parametrizations of $C^{*}$-algebras
Kechris [Kechrisdescriptivecstar] introduced a standard Borel space parametrization of the space of separable $C^{*}$-algebras. Farah, Toms and Törnquist [farah2013descriptive, Farahturbulence2014] considered a few other such parametrizations, compared them, and studied Borel complexity of several equivalence relations and invariants appearing in this context. Furthermore, in [farah2021MTofC*] criteria on Borelness of some classes of $C^{*}$-algebras are given. We survey these results and introduce some notations below.
We follow the presentation of [Farahturbulence2014, Section 2.1] for the standard Borel structure associated to a given $C^{*}$-algebra. In particular, let $H$ be a separable infinite dimensional Hilbert space and let $\mathcal{B}(H)$ be the space of bounded operators on $H$; note that $\mathcal{B}(H)$ is a standard Borel space when equipped with the Borel structure generated by the weakly open subsets. For every $\gamma\in\mathcal{B}(H)^{\mathbb{N}}$, we can associate $C^{*}(\gamma)$, the $C^{*}$-algebra generated by the sequence $\gamma$. Then since all separable $C^{*}$-algebras arise as $C^{*}$-subalgebras of $\mathcal{B}(H)$, the space $\Gamma:=\mathcal{B}(H)^{\mathbb{N}}$ gives a standard Borel parametrization for all separable $C^{*}$-algebras. We denote by $\Gamma_{u}$ the subset consisting of parameters for unital separable $C^{*}$-algebras.
On the other hand, let $\widehat{\Xi}\subset\omega^{\omega^{2}}\times\omega^{\omega^{2}}\times\omega^{\omega}\times\mathbb{R}_{\geq 0}^{\omega}\times(\omega^{\omega})^{\mathbb{Q}(i)}$ be the subset of tuples $A=(+,\times,(-)^{*},\|-\|,(\lambda\cdot)_{\lambda\in\mathbb{Q}(i)})$ that define a $C^{*}$-algebra. For the details see [Farahturbulence2014, Section 2.4]. Each $A\in\widehat{\Xi}$ defines a $C^{*}$-algebra $\widehat{A}$ which is the completion of $\omega$ with respect to the norm $\|$-$\|$. Note that $\widehat{\Xi}$ is a Borel subset of the ambient product, so it is a standard Borel space. Farah, Toms, and Törnquist prove that there is a Borel isomorphism $f\colon\Gamma\setminus\{0\}\to\widehat{\Xi}$ such that for every $\gamma\in\Gamma\setminus\{0\}$, the $C^{*}$-algebra $C^{*}(\gamma)$ is isomorphic to $\widehat{A}$ for $A=f(\gamma)$. Thus, any of these standard Borel spaces can be seen as parametrizing non-trivial separable $C^{*}$-algebras. These parametrizations are equivalent to the one we discussed in Section 2.5 in the following sense.
@@
Let us a take a brief moment to consider the assumptions of this statement. This class of $C^{*}$-algebras was briefly discussed in the introduction to this paper; each assumption here is necessary for classification, [^8] The open question asking if the UCT holds for all nuclear $C^{*}$-algebras could render the UCT assumption redundant. and a deeper discussion of each of the assumptions can be found in the introduction of [CGSTW-new-classificationI-2023]. For our purposes, recall that our Borel parametrizations are only for separable $C^{*}$-algebras. Simplicity, nuclearity, and $\mathcal{Z}$-stability were previously studied from a descriptive set-theoretic viewpoint. The presence of $\mathcal{Z}$-stability is Borel computable by [farah2013descriptive, Theorem 1.1(vi)], while simplicity and nuclearity are model-theoretically accessible: they are definable by uniform families of formulas [farah2021MTofC*, Theorem 5.7.3(3), (6)], and therefore Borel-computable by [farah2021MTofC*, Proposition 5.17.1]. This establishes the following.
@@
We now turn our attention to the assumption of Rosenberg and Schochet’s UCT [RosenbergSchochetUCT]. In their original formulation, the UCT relates the Kasparov group $KK(A,B)$ with the $K$-theory of $A$ and $B$. For the purposes of this paper, we only require the following fact from [RosenbergSchochetUCT, Section 7].
@@
As such we now turn our attention to $KK$-equivalence (that we introduce in Definition 3.7). For separable $C^{*}$-algebras A,BA,\ B, the abelian groups $KK(A,B)$ were originally defined by G.G. Kasparov in [Kasparov1981-origKK]; these can be vaguely thought of as a group of generalized homomorphisms between $C^{*}$-algebras. These $KK$-groups have since been studied from a number of viewpoints, but we only require the category-theoretic characterization obtained by Higson in [Higson-KKtheory].
@@
Theorem 3.6 (Theorem 4.5 in [Higson-KKtheory]).
@@ 3.2. The invariant classifying $C^{*}$-algebras
For a unital separable $C^{*}$-algebra $A$, the invariant $KT_{u}(A)$ given in the unital classification theorem consists of unordered operator-algebraic $K$-theory, along with the space of all traces on $A$ and a pairing map which relates traces and $K$-theory. We set out a quick description of the components of the invariant to formalize $KT_{u}(\cdot)$ in first-order logic; those interested in a more thorough introduction of it, see [CGSTW-new-classificationI-2023, §2.1].
@@
[CGSTW-new-classificationI-2023, Definition 2.3] $KT_{u}$ is the functor on unital separable $C^{*}$-algebras which assigns to $A$ the quadruple
@@
Theorem 3.10 (Theorem 3.3, [farah2013descriptive]).
@@
Since this is only a translation of results from [farah2013descriptive, Farahturbulence2014] to our setting, in the following proof we freely use notations and definitions from loc.cit.
@@
By Lemma 3.1 and [Farahturbulence2014, Lemma 3.15], it is enough to prove that there is a Borel map $\mathbf{Ell}\rightarrow\operatorname{Mod}_{\omega}(L_{\operatorname{inv}})$ that sends
@@
Now we argue that there is a Borel map $\delta\colon\mathbf{K}_{\textnormal{conv}}\to\operatorname{Mod}_{\omega}(L_{ous})$ sending $T$ to a code for the separable $L_{ous}$-structure coming from the Archimedean order unit space $\operatorname{Aff}(T)$. By [Farahturbulence2014, Lemma 4.4] there is a Borel map $\Psi\colon\mathbf{K}_{\textnormal{conv}}\to S\bigl(C(\Delta,\mathbb{R})\bigr)$ that maps a compact convex set $K\subseteq\Delta^{\mathbb{N}}$ to a closed subspace $\Psi(K)$ of the Banach algebra $C(\Delta,\mathbb{R})$ (equipped with the supremum norm), that is isomorphic to $\operatorname{Aff}(K)$. By the Kuratowski–Ryll-Nardzewski measurable selection theorem, there are Borel maps
@@
The last thing to check is that computing the pairing $\rho$ from a tuple $(G_{0},G_{1},T,r)$ is Borel. Recall that the map $\Psi\colon\mathbf{K}_{\textnormal{conv}}\to S\bigl(C(\Delta,\mathbb{R})\bigr)$ from [Farahturbulence2014, Lemma 4.4] is constructed as follows:
  • a continuous map $\chi^{\prime}\colon\mathbf{K}_{\textnormal{conv}}\to C(\Delta^{\mathbb{N}},\Delta^{\mathbb{N}})$ is constructed (note that in [farah2013descriptive, Lemma 3.1] this map is called $\Psi$), such that $\chi^{\prime}(K)$ is a continuous retraction onto $K$;
@@
By [farah2013descriptive, Definition 3.2], $r\in\operatorname{Pairing}(T,G_{0})$, which means that $r=r^{\prime}\circ\chi^{\prime}(T)$, for some $r^{\prime}\colon T\to\operatorname{States}(G_{0})\subseteq\mathbb{R}^{n}$ (where $n$ corresponds to the cardinality of the group encoded by $G_{0}$). We define $\rho=\rho_{(G_{0},G_{1},T,r)}\colon n\to\Psi(T)\subseteq C(\Delta,\mathbb{R})$ by
@@ 3.3. Borel properties of separable $C^{*}$-algebras
The above lemma is related to [Farahturbulence2014, Lemma 3.10], where the Borelness of $G_{0}$ (and of the functor of tensoring with any separable nuclear $C^{*}$-algebra) is shown. We conclude with the analyticity of the set of pairs of homotopic morphisms.
@@ 3.5. Questions
If the UCT restricted to the Borel set of nuclear $C^{*}$-algebras results in a strictly analytic set, then this would answer the UCT problem in the negative, as the set of nuclear $C^{*}$-algebras would differ from the set of nuclear $C^{*}$-algebras with the UCT. A relaxation of the above question would be to ask the same for the $KK$-equivalence relation (see also [Farahturbulence2014, Problem 9.3]).
@@
The following is also natural to ask, and is related to Problems 9.4, 9.7 in [Farahturbulence2014].
@@
Note that in Corollary 3.11 we only prove Borelness on the level of objects. Borelness of the functor would imply that it comes from an infinitary interpretation, see [MR3893282, MR4029718] (one would need an appropriate continuous logic generalisation).

### v2

@@ 2.5. Borel category of $L$-structures
The map $\pi$ is precisely the quotient map under the equivalence $(A,B,f)\sim(A,B,g)\iff\widehat{f}=\widehat{g}$, and the above lemma verifies that the quotient is itself a standard Borel space, and $\pi$ is a Borel map. In the language of Borel equivalence relations, this means that $\sim$ is smooth (see for example [41, Definition 3.11]). We have successfully isolated a standard Borel space – namely $\pi(X_{1})$ – that is in natural bijection with the family of morphisms in $\mathcal{C}_{T}$, and will serve as the space of morphisms in the Borel category $\mathbf{X}^{T}$. Thus, we write $X^{T}_{\operatorname{Mor}}:=\pi(X_{1})$. We define $s,t\colon X^{T}_{\operatorname{Mor}}\to X^{T}_{\operatorname{Ob}}$ be the first and second projections respectively, and let $i\colon X^{T}_{\operatorname{Ob}}\to X^{T}_{\operatorname{Mor}}$ be the map sending $A$ to $\pi(A,A,i_{A})$, where $i_{A}(a,b):=d_{A}(a,b)$. We define $\circ\colon X^{T}_{\operatorname{Mor}}\times_{t,X^{T}_{\operatorname{Ob}},s}X^{T}_{\operatorname{Mor}}\to X^{T}_{\operatorname{Mor}}$ to come from composition of homomorphisms in $\mathcal{C}_{T}$.
@@
$\operatorname{\bf{Mod}}_{\omega}(T)$ is a Borel model of $\mathcal{C}_{T}$.
@@
Let $\widehat{\,\cdot\,}\colon\operatorname{Mod}_{\omega}(T)\to\mathcal{C}_{T}$ take any $A\in X^{T}_{\operatorname{Ob}}$ to $\widehat{A}$ and any $(A,B,\delta)\in X^{T}_{\operatorname{Mor}}(A,B)$ to the unique L{L}-homomorphism $\widehat{f}\colon\widehat{A}\to\widehat{B}$ with $\delta(a,b)=d_{B}(\widehat{f}(a),b)$ for each $a\in A$ and $b\in B$. We have constructed $s,t,i$ and $\circ$ precisely such that this is a functor. Moreover, for fixed $A,B\in X^{T}_{\operatorname{Ob}}$, $\widehat{\,\cdot\,}$ is bijective from $X^{T}_{\operatorname{Mor}}(A,B)$ to the set of $L$-homomorphisms by construction (see Lemma 2.37). In other words, $\widehat{\,\cdot\,}$ is fully faithful.
@@ 3. Application to $C^{*}$-algebras
The desire to classify objects is ubiquitous across mathematical disciplines; here we discuss one particular classification program for norm-closed subalgebras of bounded operators of Hilbert spaces, i.e., $C^{*}$-algebras. See [5, § 1.1, § 1.2] for a survey and brief history of this topic. We assume familiarity with $C^{*}$-algebras, whose model theory is laid out in [15, § 2.3.1], see also [14]. We start by presenting the unital classification theorem along with the classifying invariant, in order to set out a first-order language suitable to understanding the invariant.
@@ 3.1. Borel parametrizations of $C^{*}$-algebras
Kechris [40] introduced a standard Borel space parametrization of the space of separable $C^{*}$-algebras. Farah, Toms and Törnquist [16, 17] considered a few other such parametrizations, compared them, and studied Borel complexity of several equivalence relations and invariants appearing in this context. Furthermore, in [14] criteria on Borelness of some classes of $C^{*}$-algebras are given. We survey these results and introduce some notations below.
We follow the presentation of [17, Section 2.1] for the standard Borel structure associated to a given $C^{*}$-algebra. In particular, let $H$ be a separable infinite dimensional Hilbert space and let $\mathcal{B}(H)$ be the space of bounded operators on $H$; note that $\mathcal{B}(H)$ is a standard Borel space when equipped with the Borel structure generated by the weakly open subsets. For every $\gamma\in\mathcal{B}(H)^{\mathbb{N}}$, we can associate $C^{*}(\gamma)$, the $C^{*}$-algebra generated by the sequence $\gamma$. Then since all separable $C^{*}$-algebras arise as $C^{*}$-subalgebras of $\mathcal{B}(H)$, the space $\Gamma:=\mathcal{B}(H)^{\mathbb{N}}$ gives a standard Borel parametrization for all separable $C^{*}$-algebras. We denote by $\Gamma_{u}$ the subset consisting of parameters for unital separable $C^{*}$-algebras.
On the other hand, let $\widehat{\Xi}\subset\omega^{\omega^{2}}\times\omega^{\omega^{2}}\times\omega^{\omega}\times\mathbb{R}_{\geq 0}^{\omega}\times(\omega^{\omega})^{\mathbb{Q}(i)}$ be the subset of tuples $A=(+,\times,(-)^{*},\|-\|,(\lambda\cdot)_{\lambda\in\mathbb{Q}(i)})$ that define a $C^{*}$-algebra. For the details see [17, Section 2.4]. Each $A\in\widehat{\Xi}$ defines a $C^{*}$-algebra $\widehat{A}$ which is the completion of $\omega$ with respect to the norm $\|$-$\|$. Note that $\widehat{\Xi}$ is a Borel subset of the ambient product, so it is a standard Borel space. Farah, Toms, and Törnquist prove that there is a Borel isomorphism $f\colon\Gamma\setminus\{0\}\to\widehat{\Xi}$ such that for every $\gamma\in\Gamma\setminus\{0\}$, the $C^{*}$-algebra $C^{*}(\gamma)$ is isomorphic to $\widehat{A}$ for $A=f(\gamma)$. Thus, any of these standard Borel spaces can be seen as parametrizing non-trivial separable $C^{*}$-algebras. These parametrizations are equivalent to the one we discussed in Section 2.5 in the following sense.
@@
Let us a take a brief moment to consider the assumptions of this statement. This class of $C^{*}$-algebras was briefly discussed in the introduction to this paper; each assumption here is necessary for classification, [^8] The open question asking if the UCT holds for all nuclear $C^{*}$-algebras could render the UCT assumption redundant. and a deeper discussion of each of the assumptions can be found in the introduction of [5]. For our purposes, recall that our Borel parametrizations are only for separable $C^{*}$-algebras. Simplicity, nuclearity, and $\mathcal{Z}$-stability were previously studied from a descriptive set-theoretic viewpoint. The presence of $\mathcal{Z}$-stability is Borel computable by [16, Theorem 1.1(vi)], while simplicity and nuclearity are model-theoretically accessible: they are definable by uniform families of formulas [14, Theorem 5.7.3(3), (6)], and therefore Borel-computable by [14, Proposition 5.17.1]. This establishes the following.
@@
We now turn our attention to the assumption of Rosenberg and Schochet’s UCT [49]. In their original formulation, the UCT relates the Kasparov group $KK(A,B)$ with the $K$-theory of $A$ and $B$. For the purposes of this paper, we only require the following fact from [49, Section 7].
@@
As such we now turn our attention to $KK$-equivalence (that we introduce in Definition 3.7). For separable $C^{*}$-algebras A,BA,\ B, the abelian groups $KK(A,B)$ were originally defined by G.G. Kasparov in [38]; these can be vaguely thought of as a group of generalized homomorphisms between $C^{*}$-algebras. These $KK$-groups have since been studied from a number of viewpoints, but we only require the category-theoretic characterization obtained by Higson in [31].
@@
Theorem 3.6 (Theorem 4.5 in [31]).
@@ 3.2. The invariant classifying $C^{*}$-algebras
For a unital separable $C^{*}$-algebra $A$, the invariant $KT_{u}(A)$ given in the unital classification theorem consists of unordered operator-algebraic $K$-theory, along with the space of all traces on $A$ and a pairing map which relates traces and $K$-theory. We set out a quick description of the components of the invariant to formalize $KT_{u}(\cdot)$ in first-order logic; those interested in a more thorough introduction of it, see [5, §2.1].
@@
[5, Definition 2.3] $KT_{u}$ is the functor on unital separable $C^{*}$-algebras which assigns to $A$ the quadruple
@@
Theorem 3.10 (Theorem 3.3, [16]).
@@
Since this is only a translation of results from [16, 17] to our setting, in the following proof we freely use notations and definitions from loc.cit.
@@
By Lemma 3.1 and [17, Lemma 3.15], it is enough to prove that there is a Borel map $\mathbf{Ell}\rightarrow\operatorname{Mod}_{\omega}(L_{\operatorname{inv}})$ that sends
@@
Now we argue that there is a Borel map $\delta\colon\mathbf{K}_{\textnormal{conv}}\to\operatorname{Mod}_{\omega}(L_{ous})$ sending $T$ to a code for the separable $L_{ous}$-structure coming from the Archimedean order unit space $\operatorname{Aff}(T)$. By [17, Lemma 4.4] there is a Borel map $\Psi\colon\mathbf{K}_{\textnormal{conv}}\to S\bigl(C(\Delta,\mathbb{R})\bigr)$ that maps a compact convex set $K\subseteq\Delta^{\mathbb{N}}$ to a closed subspace $\Psi(K)$ of the Banach algebra $C(\Delta,\mathbb{R})$ (equipped with the supremum norm), that is isomorphic to $\operatorname{Aff}(K)$. By the Kuratowski–Ryll-Nardzewski measurable selection theorem, there are Borel maps
@@
The last thing to check is that computing the pairing $\rho$ from a tuple $(G_{0},G_{1},T,r)$ is Borel. Recall that the map $\Psi\colon\mathbf{K}_{\textnormal{conv}}\to S\bigl(C(\Delta,\mathbb{R})\bigr)$ from [17, Lemma 4.4] is constructed as follows:
  • a continuous map $\chi^{\prime}\colon\mathbf{K}_{\textnormal{conv}}\to C(\Delta^{\mathbb{N}},\Delta^{\mathbb{N}})$ is constructed (note that in [16, Lemma 3.1] this map is called $\Psi$), such that $\chi^{\prime}(K)$ is a continuous retraction onto $K$;
@@
By [16, Definition 3.2], $r\in\operatorname{Pairing}(T,G_{0})$, which means that $r=r^{\prime}\circ\chi^{\prime}(T)$, for some $r^{\prime}\colon T\to\operatorname{States}(G_{0})\subseteq\mathbb{R}^{n}$ (where $n$ corresponds to the cardinality of the group encoded by $G_{0}$). We define $\rho=\rho_{(G_{0},G_{1},T,r)}\colon n\to\Psi(T)\subseteq C(\Delta,\mathbb{R})$ by
@@ 3.3. Borel properties of separable $C^{*}$-algebras
The above lemma is related to [17, Lemma 3.10], where the Borelness of $G_{0}$ (and of the functor of tensoring with any separable nuclear $C^{*}$-algebra) is shown. We conclude with the analyticity of the set of pairs of homotopic morphisms.
@@ 3.5. Questions
If the UCT restricted to the Borel set of nuclear $C^{*}$-algebras results in a strictly analytic set, then this would answer the UCT problem in the negative, as the set of nuclear $C^{*}$-algebras would differ from the set of nuclear $C^{*}$-algebras with the UCT. A relaxation of the above question would be to ask the same for the $KK$-equivalence relation (see also [17, Problem 9.3]).
@@
The following is also natural to ask, and is related to Problems 9.4, 9.7 in [17].
@@
Note that in Corollary 3.11 we only prove Borelness on the level of objects. Borelness of the functor would imply that it comes from an infinitary interpretation, see [28, 7] (one would need an appropriate continuous logic generalisation).